d²y/dx²-2tanx=0, dy/DX+5y=e^xsecx


1
Expert's answer
2020-03-31T16:07:31-0400

"1)\frac{d^2y}{dx^2}-2\tan x=0\\
y''-2\tan x=0\\
 y''=2\tan x\\
\frac{dy'}{dx}=2\frac{\sin x}{\cos x}\\
y'=2\int\frac{\sin x}{\cos x}dx\\
y'=-2ln|\cos x|+c_1\\
\frac{dy}{dx}=-2ln|\cos x|+c_1\\
y=\int(-2ln|\cos x|+c_1)dx\\
y=-2xln|\cos x|-2\int x\tan x dx+xc_1+c_2"


"2) \frac{dy}{dx}+5y=\frac{e^x}{\cos x}\\
y'+5y=\frac{e^x}{\cos x}\\
i) y'+5y=0\\
\frac{dy}{dx}=-5y\\
\frac{dy}{y}=-5dx\\
ln|y|=-5x+ln|c|\\
y_0=c\cdot e^{-5x}\\
ii) y=c(x)\cdot e^{-5x}\\
c'e^{-5x}-5ce^{-5x}+5ce^{-5x}=\frac{e^x}{\cos x}\\
c'=\frac{e^{6x}}{\cos x}\\
c=\int \frac{e^{6x}}{\cos x}dx+c_1\\
y=(\int \frac{e^{6x}}{\cos x}dx+c_1)\cdot e^{-5x}"